# Vision-Language-Action Models: Concepts, Progress, Applications and Challenges

### v1

@@ 3 Progress in Vision-Language-Action Models
The inception of VLA models was catalyzed by the remarkable success of transformer-based LLMs, notably ChatGPT, released in November 2022, which demonstrated unprecedented semantic reasoning capabilities (ChatGPT) [139]. This breakthrough inspired researchers to extend language models to multimodal domains, integrating perception and action for robotics. By 2023, GPT-4 introduced multimodal capabilities, processing both text and images, which spurred efforts to incorporate physical actions (GPT-4) [1]. Concurrently, VLMs like CLIP (2022) [157] and Flamingo (2022) [3] had established robust visual-text alignment through contrastive learning, enabling zero-shot object recognition and laying the groundwork for VLA models (CLIP). These models leveraged large-scale web datasets to align images with textual descriptions, a critical precursor to integrating actions.
A pivotal development was the creation of large-scale robotic datasets, such as RT-1’s 130,000 demonstrations, which provided action-grounding data essential for co-training vision, language, and action components [18]. These datasets captured diverse tasks and environments, enabling models to learn generalizable behaviors. Architectural breakthroughs followed with Google’s RT-2 in 2023 [17], a landmark VLA model that unified vision, language, and action tokens, treating robotic control as an autoregressive sequence prediction task (RT-2 Blog). RT-2 discretized actions using Discrete Cosine Transform (DCT) compression and Byte-Pair Encoding (BPE), achieving a 63% improvement in performance on novel objects. Multimodal fusion techniques, such as cross-attention transformers, integrated Vision Transformer (ViT)-processed images (e.g., 400 patch tokens) with language embeddings, enabling robots to execute complex commands like “Pick the red cup left of the bowl.” Additionally, UC Berkeley’s Octo model (2023) introduced an open-source approach with 93M parameters and diffusion decoders, trained on 800,000 robot demonstrations from the OpenX-Embodiment Dataset, further broadening the research landscape [167].
@@ 3.1 Architectural Innovations in VLA Models
From 2023 to 2024, VLA models underwent significant architectural advancements and refined training methodologies. Dual-system architectures emerged as a key innovation, exemplified by NVIDIA’s Groot N1 (2025) [13], which combined System 1 (fast diffusion policies with 10ms latency for low-level control) and System 2 (LLM-based planners for high-level task decomposition). This separation enabled efficient coordination between strategic planning and real-time execution, enhancing adaptability in dynamic environments. Other models, like Stanford’s OpenVLA (2024) [94], introduced a 7B-parameter open-source VLA trained on 970k real-world robot demonstrations, using dual vision encoders (DINOv2 [128] and SigLIP [204]) and a Llama 2 language model [172], outperforming larger models like RT-2-X (55B) [94]. Training paradigms evolved to leverage co-fine-tuning on web-scale vision-language data (e.g., LAION-5B) [152] and robotic trajectory data (e.g., RT-X) [175], aligning semantic knowledge with physical constraints [152]. Synthetic data generation tools like UniSim addressed data scarcity by creating photorealistic scenarios, such as occluded objects, crucial for robust training (UniSim [200]). Parameter efficiency was enhanced through Low-Rank Adaptation (LoRA) adapters [72], which allowed domain adaptation without full retraining, reducing GPU hours by 70%. The introduction of diffusion-based policies, as seen in Physical Intelligence’s pi 0 model (2024) [14], offered improved action diversity but required significant computational resources. These advancements democratized VLA technology, fostering collaboration and accelerating innovation.
@@
1. Early Fusion Models: One class of approaches focuses on fusing vision and language representations at the input stage before passing them to the policy module. Huang et al.’s EF-VLA model [74], presented at ICLR 2025, exemplifies this trend by retaining the representational alignment established by CLIP [157]. EF-VLA accepts image-text pairs, encodes them with CLIP’s frozen encoders, and fuses the resulting embeddings early in the transformer backbone—prior to action prediction. This design ensures that the semantic consistency learned during CLIP pretraining is preserved, reducing overfitting and enhancing generalization. Notably, EF-VLA demonstrated a 20% performance improvement on compositional manipulation tasks and reached 85% success on previously unseen goal descriptions. By avoiding fine-tuning of the vision-language modules, this approach also preserves computational efficiency and prevents catastrophic forgetting during domain-specific training.
@@
2. Dual-System Architectures: Inspired by dual-process theories of human cognition, models like NVIDIA’s Groot N1 (2025) [13] implement two complementary subsystems: a fast-reactive module (System 1) and a slow-reasoning planner (System 2). System 1 comprises a diffusion-based control policy that operates at 10 ms latency, ideal for fine-grained, low-level control such as end-effector stabilization or adaptive grasping. In contrast, System 2 uses a LLM for task planning, skill composition, and high-level sequencing. The planner parses long-horizon goals (e.g., “clean the table”) into atomic subtasks, while the low-level controller ensures real-time execution. This decomposition enables multi-timescale reasoning and improved safety, especially in environments where rapid reaction and deliberation must co-exist. In benchmark tests on multi-stage household manipulation, Groot N1 outperformed monolithic models by 17% in success rate and reduced collision failures by 28
@@
3. Self-Correcting Frameworks: A third architectural evolution is the development of self-correcting VLA models, designed to detect and recover from failure conditions without external supervision. SC-VLA (2024) introduces a hybrid execution loop featuring a fast inference path and a slow correction path. In this framework, the default behavior is to predict poses or actions directly from the fused embedding using a lightweight transformer. When failures are detected—e.g., unsuccessful grasps or obstacle collisions—the model invokes a secondary process that performs chain-of-thought reasoning [211, 203]. This path queries an internal LLM (or external expert system) to diagnose failure modes and generate correction strategies [48]. For example, if the robot repeatedly misidentifies an occluded object, the LLM may suggest an active viewpoint change or gripper reorientation. In closed-loop experiments, SC-VLA reduced task failure rates by 35% and significantly improved recoverability in cluttered and adversarial environments.
@@
VLA models exhibit a rich diversity of architectural designs and functional emphases, which can be systematically organized along the dimensions of end-to-end versus modular pipelines, hierarchical versus flat policy structures, and the balance between low-level control and high-level planning (Table 1). End-to-end VLAs, such as CLIPort [157], RT-1 [18], and OpenVLA [94], process raw sensory inputs directly into motor commands via a single unified network. By contrast, component-focused models like VLATest [182] and Chain-of-Affordance [100] decouple perception, language grounding, and action modules, enabling targeted improvements in individual submodules.
@@
Table 1 further highlights how recent VLAs balance these trade-offs. Models such as OpenDriveVLA [220] and CombatVLA [29] prioritize hierarchical planning in dynamic, safety-critical domains, whereas lightweight, edge-targeted systems like Edge VLA [19] and TinyVLA [186] emphasize real-time low-level policies at the expense of high-level reasoning. This classification framework not only clarifies the design space of VLAs but also guides future development by pinpointing under-explored combinations—such as fully end-to-end, hierarchical models optimized for embedded deployment—that promise to advance both the capabilities and the applicability of VLA systems across robotics, autonomous driving, and beyond.
The classification in Table 1 is significant because it provides a clear framework for comparing diverse VLA architectures, highlighting how design choices—such as end-to-end integration versus hierarchical decomposition—impact task performance, scalability, and adaptability. By categorizing models along dimensions like low-level policy execution and high-level planning, researchers can pinpoint strengths and limitations of existing approaches and identify opportunities for innovation. This taxonomy aids in selecting appropriate architectures for specific applications (e.g., real‐time control vs. strategic reasoning) and guides future development toward hybrid systems that balance responsiveness with cognitive planning, ultimately accelerating progress in embodied AI. Additionally, to synthesize recent advancements in VLA models, Table 2 presents a comparative summary of notable systems developed from 2022 through 2025. Building upon architectural innovations such as early fusion, dual-system processing, and self-correcting feedback loops, these models incorporate diverse design philosophies and training strategies. Each entry highlights the model’s key components—vision and language encoders, action decoders—and the datasets used to ground their capabilities. Models like CLIPort [157] and RT-2 [224] laid early foundations by aligning semantic embeddings with action policies, while more recent frameworks like Pi-Zero, CogACT [102], and Groot N1 [13] introduce scalable architectures with diffusion-based or high-frequency controllers. Several models leverage multimodal pretraining with internet-scale vision-language corpora and robot trajectory datasets, enhancing generalization and zero-shot capabilities [223, 219, 218, 198]. This tabulated comparison serves as a reference point for researchers seeking to understand the functional diversity, domain applicability, and emerging trends in VLA design across real and simulated environments.
@@
Table 2: Summary of VLA models, detailing each model’s name, architecture features, training dataset, and highlighting their key strengths or unique capabilities in robotics and AI tasks.
Model (Reference) Architecture Components Training Dataset Key Strength / Uniqueness
CLIPort [157] 1. Vision Encoder: CLIP-ResNet50 + Transporter-ResNet 2. Language Encoder: CLIP-GPT 3. Action Decoder: LingUNet Self-collected [SC] Combines semantic CLIP features with spatial Transporter network for precise SE(2) manipulation.
RT-1 [18] 1. Vision Encoder: EfficientNet 2. Language Encoder: Universal Sentence Encoder 3. Action Decoder: Transformer RT-1-Kitchen [SC] Pioneering Transformer architecture with discretized actions for multi-task kitchen manipulation.
RT-2 [224] 1. Vision Encoder: ViT-22B/ViT-4B 2. Language Encoder: PaLI-X/PaLM-E 3. Action Decoder: Symbol-tuning VQA + RT-1-Kitchen First large VLA co-finetuned on internet VQA data and robot data for emergent capabilities.
Gato [141] 1. Vision Encoder: ViT 2. Language Encoder: SentencePiece 3. Action Decoder: Transformer Self-collected [SC] Generalist agent handling Atari, captioning, and robotics through unified tokenization.
VIMA [86] 1. Vision Encoder: ViT + Mask R-CNN 2. Language Encoder: T5 3. Action Decoder: Transformer VIMA-Data [SC] Multi-modal prompt handling with 6 types of vision-language grounding tasks.
ACT [216] 1. Vision Encoder: ResNet-18 2. Language Encoder: 3. Action Decoder: CVAE-Transformer ALOHA [SC] Temporal ensembling for smooth bimanual manipulation with 0.1mm precision.
Octo [167] 1. Vision Encoder: CNN 2. Language Encoder: T5-base 3. Action Decoder: Diffusion Transformer Open X-Embodiment First policy trained on 4M+ robot trajectories from 22 robot types.
VoxPoser [78] 1. Vision Encoder: ViLD + MDETR 2. Language Encoder: GPT-4 3. Action Decoder: MPC Zero-shot LLM+VLM composition for constraint-aware motion planning without training.
Diffusion Policy [34] 1. Vision Encoder: ResNet-18 2. Language Encoder: 3. Action Decoder: U-Net/Transformer Self-collected [SC] Pioneering diffusion-based visuomotor policy handling multimodal action distributions.
OpenVLA [94] 1. Vision Encoder: DINOv2 + SigLIP 2. Language Encoder: Prismatic-7B 3. Action Decoder: Symbol-tuning OXE + DROID Open-source alternative to RT-2 with efficient LoRA fine-tuning.
(Pi-Zero) [14] 1. Vision Encoder: PaliGemma VLM backbone 2. Language Encoder: PaliGemma (multimodal) 3. Action Decoder: 300M-parameter diffusion model Pi-Cross-Embodiment Robot dataset Lightweight, efficient VLA model ( 3B params) excelling at general robot control and bimanual manipulation, with strong open-world generalization across diverse robots and tasks [1][5][7].
Pi-0 Fast [133] 1. Vision Encoder: PaliGemma VLM backbone 2. Language Encoder: PaliGemma (multimodal) 3. Action Decoder: Autoregressive Transformer with FAST (Frequency-space Action Sequence Tokenization) Pi- Cross-Embodiment Robot dataset Variant of Pi-0 optimized for high-frequency, real-time control using compressed action tokens; achieves up to 15x faster inference for discrete robot actions and strong generalization.
@@
RDT-1B [112] 1. Vision Encoder: Multi-view RGB image encoder 2. Language Encoder: Transformer-based language module 3. Action Decoder: Diffusion Transformer with unified action space 1M+ multi-robot episodes (46 datasets), fine-tuned on 6K+ bimanual ALOHA episodes 1.2B-parameter diffusion foundation model for bimanual manipulation; excels at language-conditioned, dexterous control and zero-shot generalization, with strong but task-specific performance in multi-object settings.
Helix  [^1] https://www.figure.ai/news/helix 1. Vision Encoder: Open-source VLM (System 2) for multimodal scene and language understanding at 7–9 Hz 2. Language Encoder: Integrated with VLM for broad generalization and semantic comprehension 3. Action Decoder: Transformer-based visuomotor policy (System 1) for continuous, full upper-body control at 200 Hz End-to-end on Figure robot data (pixels and language to actions) First VLA model for real-time, high-DoF humanoid control; enables zero-shot generalization, fine-grained dexterity, and collaborative multi-robot manipulation in open-world tasks.
@@
CogACT [102] 1. Vision Encoder: DINOv2 ViT-L/14, SigLIP ViT-So400M/14 2. Language Encoder: Llama-2 (via Prismatic-7B VLM) 3. Action Decoder: Diffusion Transformer (DiT-Base, 300M parameters) Open X-Embodiment (OXE) subset, real-world Realman & Franka tasks Componentized VLA with specialized diffusion action transformer; outperforms OpenVLA by 59.1% in real-world success, excels at adaptation and generalization to new robots and unseen objects.
Chain-of-Affordance (CoA) [100] 1. Vision Encoder: Visual encoder with affordance feature extraction 2. Language Encoder: Transformer-based language module for sequential reasoning prompts 3. Action Decoder: Autoregressive and diffusion policy with affordance-conditioned outputs LIBERO benchmark, real and simulated manipulation tasks Incorporates reasoning via sequential affordances (object, grasp, spatial, movement); achieves superior LIBERO performance over OpenVLA, excelling in spatial reasoning and obstacle avoidance for precise task completion.
@@
Edge VLA (EVLA) [19] 1. Vision Encoder: SigLIP + DINOv2 2. Language Encoder: Qwen2 (0.5B parameters) 3. Action Decoder: Joint control prediction (non-autoregressive) Bridge dataset, OXE, 1.2M text-image pairs Lightweight VLA model optimized for edge devices (e.g., Jetson Nano) with 30–50 Hz inference; achieves performance comparable to OpenVLA while enabling efficient, real-time deployment on low-power hardware.
@@
ShowUI-2B [108] 1. Vision Encoder: UI-Guided Visual Token Selection (transformer-based) 2. Language Encoder: Interleaved vision-language-action streaming 3. Action Decoder: Transformer for GUI action sequence prediction 256K high-quality GUI instruction-following dataset Lightweight 2B-parameter VLA specialized for digital task automation; excels at GUI/web navigation and screenshot grounding with efficient token selection and unified vision-language-action reasoning.
@@
Groot N1 [13] 1. Vision Encoder: NVIDIA Eagle-2 VLM backbone (vision-language) 2. Language Encoder: Integrated with VLM for high-level planning and reasoning 3. Action Decoder: Diffusion Transformer (DiT) for precise, high-frequency action generation Multimodal data: human demonstrations, robot trajectories, synthetic simulation, and internet video Hybrid dual-system architecture for generalist humanoid robots, combining high-level planning with diffusion-based execution; enables dexterous, multi-step control and strong generalization across tasks and embodiments.
@@
Seer [63] 1. Vision Encoder: Visual backbone optimized for grounding and perception 2. Language Encoder: Transformer-based language module 3. Action Decoder: Autoregressive action prediction head LIBERO benchmark VLA model focused on visual perception and action prediction; achieves competitive results with strong visual grounding for manipulation, but is outperformed by OpenVLA-OFT [1][6][8][10].
@@
DiffusionVLA [184] 1. Vision Encoder: Transformer-based visual encoder for contextual perception 2. Language Encoder: Autoregressive reasoning module with next-token prediction 3. Action Decoder: Diffusion policy head for robust action sequence generation LIBERO benchmark, factory sorting, zero-shot bin-picking tasks Leverages diffusion-based action modeling for precise control; demonstrates robustness and interpretability, but is less generalizable than CoA in spatial configurations.
@@
NaVILA [32] 1. Vision Encoder: CLIP + CNN 2. Language Encoder: LLaMA-2 (task command + nav goal) 3. Action Decoder: Two-level controller: topological graph planner + RL-based locomotion Real-world legged robot nav demos Modular hierarchy enables robust terrain generalization and 88% real-world nav success using natural language
@@
RoboNurse-VLA [103] 1. Vision Encoder: SAM2 + RGB-D 2. Language Encoder: LLaMA 2 + voice-to-text encoder 3. Action Decoder: Joint pose regression with gripper classifier Surgical handover videos and voice prompts Enables accurate, real-time surgical tool handover; strong robustness to tool novelty and dynamic OR scenes
@@
Mobility VLA [35] 1. Vision Encoder: Long-context ViT + goal image encoder 2. Language Encoder: T5-based instruction encoder 3. Action Decoder: Graph planner with visual goal localization MINT dataset: vision-language instruction tours Robust navigation from multimodal input; generalizes across large unseen spaces via topological mapping
TinyVLA [186] 1. Vision Encoder: FastViT with low-latency encoding 2. Language Encoder: Compact language encoder (128-d) 3. Action Decoder: Diffusion policy decoder (50M params) Mini-ALOHA + SC tasks Outperforms OpenVLA in speed and precision; does not require pretraining; inference 5x faster with minimal compute
@@
QUAR-VLA [43] 1. Vision Encoder: CLIP + proprioceptive embedding 2. Language Encoder: BERT + custom grounding adapter 3. Action Decoder: Transformer for full-body command decoding QUART dataset (locomotion + manipulation) Quadruped-specific control with strong sim-to-real transfer and fine-grained instruction alignment
@@
ChatVLA [221] 1. Vision Encoder: Vision encoder integrated with Phase-Aligned transformer 2. Language Encoder: Prismatic LLM with Mixture-of-Experts 3. Action Decoder: Unified vision-language-action planner Unified chat-action dataset (web, robot) Excels at joint VQA and planning; mitigates forgetting; efficient across manipulation and conversational tasks
@@
PointVLA [96] 1. Vision Encoder: CLIP + 3D Point Cloud (via skip blocks) 2. Language Encoder: LLaMA-2 3. Action Decoder: Transformer with spatial token fusion Few-shot spatial tasks (real + sim) Excels at long-horizon and spatial reasoning tasks; avoids retraining by preserving pretrained 2D knowledge
@@
VLA-Cache [195] 1. Vision Encoder: SigLIP with token memory buffer 2. Language Encoder: Prismatic-7B 3. Action Decoder: Transformer with dynamic token reuse ALOHA + real-world sim fusion 40–50% faster inference with near-zero loss; dynamically reuses static features for real-time robotics

### v2

@@ 3 Progress in Vision-Language-Action Models
The inception of VLA models was catalyzed by the remarkable success of transformer-based LLMs, notably ChatGPT, released in November 2022, which demonstrated unprecedented semantic reasoning capabilities (ChatGPT) [179]. This breakthrough inspired researchers to extend language models to multimodal domains, integrating perception and action for robotics. By 2023, GPT-4 introduced multimodal capabilities by processing both text and images, which spurred subsequent efforts to extend language-centric multimodal foundation models toward incorporating physical action representations and control interfaces [1]. Concurrently, VLMs like CLIP (2022) [202] and Flamingo (2022) [3] had established robust visual-text alignment through contrastive learning, enabling zero-shot object recognition and laying the groundwork for VLM models (such as CLIP). These models leveraged large-scale labeled datasets to align images with textual descriptions, a critical precursor to integrating actions.
A pivotal development was the creation of large-scale robotic datasets, such as RT-1’s 130,000 demonstrations, which provided action-grounding data essential for co-training vision, language, and action components [19]. These datasets captured diverse tasks and environments, enabling models to learn generalizable behaviors. Architectural breakthroughs followed with Google’s RT-2 in 2023 [18], a landmark VLA model that unified vision, language, and action tokens, treating robotic control as an autoregressive sequence prediction task (RT-2 Blog(Source Link)). RT-2 discretized actions using Discrete Cosine Transform (DCT) compression and Byte-Pair Encoding (BPE), achieving a 63% improvement in performance on novel objects. Multimodal fusion techniques, such as cross-attention transformers, integrated Vision Transformer (ViT)-processed images (e.g., 400 patch tokens) with language embeddings, enabling robots to execute complex commands like “Pick the red cup left of the bowl.” Additionally, UC Berkeley’s Octo model (2023) introduced an open-source approach with 93M parameters and diffusion decoders, trained on 800,000 robot demonstrations from the OpenX-Embodiment dataset, further broadening the research landscape [218].
@@ 3.1 Architectural Innovations in VLA Models
From 2023 to 2024, VLA models underwent significant architectural advancements and refined training methodologies. Dual-system architectures emerged as a key innovation, exemplified by NVIDIA’s GR00T N1 (2025) [14], which combined System 1 (fast diffusion policies with 10ms latency for low-level control) and System 2 (LLM-based planners for high-level task decomposition). This separation enabled efficient coordination between strategic planning and real-time execution, enhancing adaptability in dynamic environments. Other models, like Stanford’s OpenVLA (2024) [122], introduced a 7B-parameter open-source VLA trained on 970k real-world robot demonstrations, using dual vision encoders (DINOv2 [164] and SigLIP [271]) and a Llama 2 language model [223], outperforming larger models like RT-2-X (55B) [122]. Training paradigms evolved to leverage co-fine-tuning on web-scale vision-language data (e.g., LAION-5B) [194] and robotic trajectory data (e.g., RT-X) [227], aligning semantic knowledge with physical constraints [194]. Synthetic data generation tools like UniSim addressed data scarcity by creating photorealistic scenarios, such as occluded objects, crucial for robust training (UniSim [264]). Parameter efficiency was enhanced through Low-Rank Adaptation (LoRA) adapters for fine-tuning [93], which allowed domain adaptation without full retraining, reducing GPU hours by 70%. The introduction of diffusion-based policies, as seen in Physical Intelligence’s pi 0 model (2024) [15], offered improved action diversity but required significant computational resources. These advancements democratized VLA technology, fostering collaboration and accelerating innovation.
@@
1. Early Fusion Models: One class of VLA approaches focuses on fusing vision and language representations at the input stage before passing them to the policy module. Huang et al.’s EF-VLA model [96], presented at International Conference on Learning Representations (ICLR 2025), exemplifies this trend by retaining the representational alignment established by CLIP [202]. EF-VLA accepts image-text pairs, encodes them with CLIP’s frozen encoders, and fuses the resulting embeddings early in the transformer backbone prior to action prediction. This design ensures that the semantic consistency learned during CLIP pretraining is preserved, reducing overfitting and enhancing generalization. Notably, EF-VLA demonstrated a 20% performance improvement on compositional manipulation tasks and reached 85% success on previously unseen goal descriptions. By keeping the vision-language backbone frozen, this approach preserves computational efficiency and avoids catastrophic forgetting, while domain-specific training is confined to lightweight policy or action modules, enabling task adaptation without sacrificing the model’s general-purpose visual-semantic representations.
@@
2. Dual-System Architectures: Inspired by dual-process theories of human cognition, models like NVIDIA’s GR00T N1 (2025) [14] implement two complementary subsystems: a fast-reactive module (System 1) and a slow-reasoning planner (System 2). System 1 comprises a diffusion-based control policy that operates at 10 ms latency, ideal for fine-grained, low-level control such as end-effector stabilization or adaptive grasping. In contrast, System 2 uses a LLM for task planning, skill composition, and high-level sequencing. The planner parses long-horizon goals (e.g., “clean the table”) into atomic subtasks, while the low-level controller ensures real-time execution. This decomposition enables multi-timescale reasoning and improved safety, especially in environments where rapid reaction and deliberation must co-exist. In benchmark tests on multi-stage household manipulation, GR00T N1 outperformed monolithic models such as RT-1, RT-2, and OpenVLA by 17% in success rate and reduced collision failures by 28%.
@@
3. Self-Correcting Frameworks: A third architectural evolution is the emergence of self-correcting VLA models, which augment conventional inference pipelines with explicit failure detection and recovery mechanisms. SC-VLA (2024) retains a standard fast inference path similar to earlier end-to-end or hierarchical VLA designs, but introduces an additional, slower correction path that is selectively activated to re-evaluate decisions and generate recovery actions when execution failures or inconsistencies are detected. In this framework, the default behavior is to predict poses or actions directly from the fused embedding using a lightweight transformer. When failures are detected e.g., unsuccessful grasps or obstacle collisions, the model invokes a secondary process that performs chain-of-thought reasoning [281, 270]. This path queries an internal LLM (or external expert system) to diagnose failure modes and generate correction strategies [59]. For example, if the robot repeatedly misidentifies an occluded object, the LLM may suggest an active viewpoint change or gripper reorientation. In closed-loop experiments, SC-VLA reduced task failure rates by 35% and significantly improved recoverability in cluttered and adversarial environments.
@@
4. Architectural Design Space of VLA Models VLA models exhibit a rich diversity of architectural designs and functional emphases, which can be systematically organized along the dimensions of end-to-end versus modular pipelines, hierarchical versus flat policy structures, and the balance between low-level control and high-level planning (Table 1). End-to-end VLAs, such as CLIPort [202], RT-1 [19], and OpenVLA [122], process raw sensory inputs directly into motor commands via a single unified network. By contrast, component-focused models like VLATest [237] and Chain-of-Affordance [129] decouple perception, language grounding, and action modules, enabling targeted improvements in individual submodules.
@@
Table 1 further highlights how recent VLAs balance these trade-offs. Models such as OpenDriveVLA [293] and CombatVLA [34] prioritize hierarchical planning in dynamic, safety-critical domains, whereas lightweight, edge-targeted systems like Edge VLA [20] and TinyVLA [242] emphasize real-time low-level policies at the expense of high-level reasoning. This classification framework not only clarifies the design space of VLAs but also guides future development by pinpointing under-explored combinations such as fully end-to-end, hierarchical models optimized for embedded deployment that promise to advance both the capabilities and the applicability of VLA systems across robotics, autonomous driving, and beyond.
The classification in Table 1 is significant because it provides a clear framework for comparing diverse VLA architectures, highlighting how design choices such as end-to-end integration versus hierarchical decomposition impact task performance, scalability, and adaptability. By categorizing models along dimensions such as low-level policy execution and high-level planning, researchers can more clearly identify the strengths and limitations of existing approaches and uncover opportunities for architectural innovation. For example, agricultural robotics tasks such as high-speed fruit harvesting or precision spraying benefit from architectures emphasizing fast, reactive low-level controllers, whereas applications like orchard navigation, multi-row coverage planning, or long-horizon crop monitoring require stronger high-level planning and reasoning capabilities. This taxonomy therefore aids in selecting appropriate VLA architectures for specific use cases and guides future development toward hybrid systems that balance responsiveness with cognitive planning, ultimately accelerating progress in embodied AI.
Additionally, to synthesize recent advancements in VLA models, Table LABEL:tab:vla_models_compact presents a summary of notable systems developed from 2022 through 2025. Building upon architectural innovations such as early fusion, dual-system processing, and self-correcting feedback loops, these models incorporate diverse design philosophy and training strategies. Each table entry explicitly enumerates the model’s architectural components—namely the vision encoder, language encoder, and action decoder together with the primary training datasets used to ground and evaluate the model’s capabilities. Models like CLIPort [202] and RT-2 [299] laid early foundations by aligning semantic embeddings with action policies, while more recent frameworks like Pi-Zero, CogACT [131], and GR00T N1 [14] introduce scalable architectures with diffusion-based or high-frequency controllers. Several models leverage multi-modal pretraining with internet-scale vision-language corpora and robot trajectory datasets, enhancing generalization and zero-shot capabilities [297, 291, 289, 258]. This tabulated comparison serves as a reference point for researchers seeking to understand the functional diversity, domain applicability, and emerging trends in VLA design across real and simulated environments.
@@
Long-VLA [64] 2025
RetoVLA [123] 2025
Vlaser [256] 2025
Discrete Diffusion VLA [138] 2025
Being-H0 [152] 2025
EgoVLA [257] 2025
StereoVLA [47] 2025
GeoVLA [212] 2025
EfficientVLA [260] 2025
Table 2: Compact summary of representative Vision–Language–Action (VLA) models. Each row reports the primary encoders/decoders, training data, and the main distinctive capability.
Model (Ref.) Architecture (vision / language / action) Training data Key strength / uniqueness
CLIPort [202] CLIP-ResNet50 + Transporter-ResNet / CLIP-GPT / LingUNet Self-collected [SC] Aligns semantic CLIP features with Transporter spatial reasoning for precise SE(2) manipulation.
RT-1 [19] EfficientNet / Universal Sentence Encoder / Transformer (discretized actions) RT-1-Kitchen [SC] Early large-scale transformer policy for multi-task kitchen manipulation with tokenized actions.
RT-2 [299] ViT-22B or ViT-4B / PaLI-X or PaLM-E / symbol-tuning (action tokens) VQA + RT-1-Kitchen Co-finetunes internet-scale VQA with robot data, yielding emergent generalization for embodied tasks.
Gato [181] ViT / SentencePiece / Transformer (unified token stream) Self-collected [SC] Generalist agent unifying robotics, language, and Atari via shared tokenization and a single transformer.
VIMA [112] ViT + Mask R-CNN / T5 / Transformer VIMA-Data [SC] Prompt-driven VL grounding across multiple compositional task types (six prompt modalities).
ACT [287] ResNet-18 / — / CVAE-Transformer ALOHA [SC] Temporal ensembling enables smooth bimanual imitation with fine control precision.
Octo [218] CNN / T5-base / Diffusion Transformer Open X-Embodiment (OXE) Large multi-robot policy trained on 4M+ trajectories spanning many robot embodiments.
VoxPoser [100] ViLD + MDETR / GPT-4 / MPC (LLM-guided planning) Zero-shot Composes LLM+VLM for constraint-aware motion planning without task-specific training.
Diffusion Policy [40] ResNet-18 / — / U-Net or Transformer diffusion Self-collected [SC] Diffusion modeling captures multimodal action distributions for robust visuomotor control.
OpenVLA [122] DINOv2 + SigLIP / Prismatic-7B / symbol-tuning OXE + DROID Open-source RT-2-like VLA; supports efficient LoRA adaptation and broad generalization.
$\pi^{0}$ (Pi-Zero) [15] PaliGemma VLM / PaliGemma (multimodal) / 300M diffusion action model Pi-Cross-Embodiment Lightweight general robot controller (reported $\sim$3B total) with strong cross-robot, open-world generalization and bimanual skills.
@@
RDT-1B [145] Multi-view RGB encoder / transformer language module / Diffusion Transformer (unified action space) 46 datasets (>>1M episodes) + ALOHA fine-tune 1.2B diffusion foundation model for dexterous bimanual manipulation with strong language conditioning and zero-shot transfer.
[…]
@@
CogACT [131] DINOv2 ViT-L/14 + SigLIP ViT-So400M/14 / Llama-2 via Prismatic-7B / DiT-Base (300M diffusion) OXE subset; Realman & Franka tasks Componentized VLA with diffusion action transformer; reported +59.1% real-world success vs. OpenVLA and strong adaptation to unseen robots/objects.
[…]
@@
Edge VLA (EVLA) [20] SigLIP + DINOv2 / Qwen2 (0.5B) / non-autoregressive joint control prediction Bridge; OXE; 1.2M text–image pairs Edge-optimized VLA (e.g., Jetson-class) with reported 30–50 Hz inference and OpenVLA-comparable performance under low power.
@@
ShowUI-2B [139] UI-guided visual token selection / interleaved V–L–A streaming / transformer GUI action predictor 256K GUI instruction-following Compact 2B VLA for digital automation; strong screenshot grounding and GUI/web navigation with efficient token selection.
@@
GR00T N1 [14] NVIDIA Eagle-2 VLM / integrated high-level planning / diffusion transformer (DiT) Human demos + robot trajectories + simulation + internet video Generalist humanoid dual-system design combining planning and diffusion execution for dexterous multi-step control and broad embodiment generalization.
@@
Seer [80] Grounding-optimized visual backbone / transformer language / autoregressive action head LIBERO Strong visual grounding for manipulation; competitive on LIBERO but typically below newer fine-tuned variants (e.g., OpenVLA-OFT).
@@
DiffusionVLA [240] Transformer visual encoder / autoregressive reasoning / diffusion action head LIBERO; factory sorting; zero-shot bin-picking Diffusion control improves robustness and interpretability; reported weaker spatial generalization than CoA in some configurations.
@@
NaVILA [38] CLIP + CNN / LLaMA-2 / hierarchical control: topological planner + RL locomotion Real-world legged robot nav demos Modular hierarchy for terrain generalization; reported 88% real-world navigation success from natural language.
@@
RoboNurse-VLA [132] SAM2 + RGB-D / LLaMA-2 + voice-to-text / pose regression + gripper classifier Surgical handover videos + voice prompts Real-time surgical tool handover robust to novel tools and dynamic operating-room scenes.
@@
Mobility VLA [42] Long-context ViT + goal-image encoder / T5-based instruction encoder / graph planner + visual goal localization MINT dataset (VL instruction tours) Topological mapping from multimodal tours enables navigation generalization across large unseen spaces.
[…]
@@
QUAR-VLA [54] CLIP + proprioceptive embedding / BERT + grounding adapter / transformer full-body decoder QUART (locomotion+manipulation) Quadruped-centric VLA with strong sim-to-real transfer and fine-grained instruction alignment.
@@
ChatVLA [295] Phase-aligned vision encoder / Prismatic MoE LLM / unified V–L–A planner Unified chat-action (web+robot) Joint VQA+planning; mitigates forgetting and supports conversational task execution with manipulation.
@@
PointVLA [125] CLIP + 3D point cloud fusion / LLaMA-2 / transformer with spatial token fusion Few-shot spatial tasks (real+sim) Improves long-horizon spatial reasoning by injecting 3D structure while preserving pretrained 2D knowledge.
@@
VLA-Cache [252] SigLIP + token memory buffer / Prismatic-7B / transformer with dynamic token reuse ALOHA + sim/real fusion Caches static visual tokens for efficiency; reported 40–50% faster inference with negligible performance loss.